What is the value of U in the equation; log9(3u+14) - log95 = log92u? I tried to work it myself and got an answer of -9, but I don't think it's right because there was a similar problem in my book but they didn't show you how to do it. Can you please walk me through it?
Answer by AnlytcPhil(1116)


log9(3u+14) - log95 = log9(2u)
Add log95 to both sides
log9(3u+14) = log9(2u) + log95
On the right use the rule of logarithms:
logBX + logBY = logB(XY)
logB(3u+14) = logB[(2u)×5]
logB(3u+14) = logB(10u)
Now we raise both sides to the 9 power, which amounts to
dropping the log9 on each side:
3u+14 = 10u
Solve that for u and get u=2.
Edwin
AnlytcPhil@aol.com